Circular No. 5712

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

SUPERNOVAE 1993C AND 1993E R. Kirshner and M. Kirshner, Center for Astrophysics, report that

spectrograms (range 400-800 nm) obtained on Feb. 18 at the Multiple Mirror Telescope confirm that SN 1993C is of type Ia, as reported on IAUC 5701. Also, SN 1993E shows the distinctive Hα emission of type-II supernovae; the FWZI for the Hα line is ∼ 11 000 km/s, and the peak wavelength is shifted to ∼ 671 nm. Further spectra and photometry may be valuable for determining a distance to this supernova by the ‘expanding photosphere method’.

1980 AA

Orbital elements by G. V. Williams from MPC 21783:

Epoch = 1993 Jan. 13.0 TT

T = 1993 Jan. 10.3273 TT  ω = 168.^{o}0395
e = 0.443895  Ω = 299.2262 }2000.0
q = 1.052284 AU  i = 4.1891

a = 1.892240 AU n^{o} = 0.3786517 P = 2.603 years

1993 TT α_{2000}  δ_{2000}  Δ  r  ϵ  β  V
Feb. 22 9^{h}02.^{m}38 + 1^{o}14.^{′}7 0.173 1.152 158.^{o}4 18.^{o}5 15.9
27 9 14.24 + 0 13.9 0.197 1.174 157.0 19.2 16.3
Mar. 4 9 24.59 − 0 26.7 0.223 1.198 155.4 20.2 16.6
9 9 33.88 − 0 53.4 0.252 1.223 153.5 21.2 17.0
14 9 42.43 − 1 10.4 0.283 1.250 151.3 22.4 17.3
19 9 50.48 − 1 21.1 0.317 1.277 148.9 23.7 17.7
24 9 58.24 − 1 28.4 0.353 1.306 146.4 25.0 18.0
29 10 05.85 − 1 34.3 0.392 1.335 143.7 26.3 18.3
Apr. 3 10 13.41 − 1 40.4 0.434 1.365 141.0 27.5 18.6

CORRIGENDA The Feb. 2 observations of 1993 BW_{3} on IAUC 5705 and of 1993 BX_{3}

on IAUC 5706 were all from plates taken by M. J. Drinkwater with the U.K. Schmidt Telescope.

1993 February 22 Daniel W. E. Green